OpenStudy (anonymous):

An airline flight attendant randomly chose 25 passengers on a transatlantic flight and asked them if they liked the new entertainment system. Twenty-one of those surveyed said yes. If there are 231 passengers in all, about how many would you expect to say that they liked the new entertainment system? A. 119 B. 147 C. 194 D. 210

jimthompson5910 (jim_thompson5910):

\[\Large \frac{\text{number who said yes}}{\text{number total}}=\frac{\text{number who said yes}}{\text{number total}}\] \[\Large \frac{21}{25}=\frac{x}{231}\] solve for x